OpenStudy (anonymous):

Guys I need help!! A woman drives 10 miles, then increases her speed by 10 miles per hour and drives another 25 miles. Find her original speed if she drives for 45 minutes.

OpenStudy (anonymous):

45

OpenStudy (kc_kennylau):

Let the original speed be s mph. 10/s + 25/(s+10) = 45/60

OpenStudy (kc_kennylau):

speed = distance / time distance = time * speed time = distance / speed

OpenStudy (gudden):

@denizen speed = distance upon time and time = distance upon speed let the original speed be x then 10/x + 25/ (x+10) = 45/60 (10 ( x+10) + 25(x) ) / x (x+10) = 45/60 now just solve it and you would get a quadratic equation which can be further solved to get the value of x

OpenStudy (anonymous):

hey @gudden thanks for the help one last question is the answer is 37.03

OpenStudy (gudden):

(10x + 100 + 25x )/ x^2 + 10x = 3 / 4 on cross multiplying we get (35x + 100 )4 = 3 ( x^2 + 10x ) 140x + 400 = 3x^2 + 30x 0 = 3x^2 -110x - 400 lets use quadratic equation d= 130 x = (110 - 130)/ 6 = - 20 / 6 = - 3.33 x = (110 + 130)/ 6 = 40 and hence speed = 40 mph..

OpenStudy (anonymous):

Thank you so much @gudden

OpenStudy (gudden):

ur welcome :D